Sue runs 8 kilometers in 45 minutes. At the same rate, how many kilometers would she run in 18 minutes.
Answer by stanbon(75887)

Sue runs 8 kilometers in 45 minutes. At the same rate, how many kilometers would she run in 18 minutes.
-------
Use a proportion:
x/18 = 8/45
x = 18(8/45)
---
x = 2(8/5) = 16/5 = 3 1/5 km
==================
Cheers,
Stan H.